# The ineffectiveness of the regularity lemma for bounded degree graphs

### v2

@@ 1. Introduction and main results
For sparse graphs, any homomorphism density is close to zero and hence such an approximation is not useful. Instead, one approximates them by neighborhood statistics, defined as follows. Suppose that $G$ is a finite graph with degrees bounded by $\Delta$. A rooted graph of radius $\leq r$ is a graph $F_{\bullet}$ with a distinguished root vertex such that all vertices of $F_{\bullet}$ are of distance $\leq r$ from the root. Rooted graphs are considered up to root-preserving isomorphism $\cong_{\bullet}$ of graphs. For a rooted graph $F_{\bullet}$ of radius $r$ and a (finite) graph $G$ define the corresponding $r$-neighborhood statistic in $G$ as
@@
which is equal to the probability that the radius $r$ ball of a uniform random vertex of $G$ is isomorphic as a rooted graph to $F_{\bullet}$.
@@
for all rooted graphs $F_{\bullet}$ of radius $r$.
@@
Bowen, Chapman, Lubotzky, and Vidick employed so-called subgroup tests to encode non-local games and then build on the ideas from the earlier deep results of Ji, Natarajan, Vidick, Wright, and Yuen [JNV+22] to deduce undecidability. We note that although the subgroup test approach might not be necessary for their original goal [Man25] it is indeed essential for our argument. Let us also mention that solely a counterexample to the Aldous–Lyons conjecture in itself would have been likely insufficient to show Theorem 3, but, the insight provided by the undecidability results suffices for this purpose.
@@
In Section 2 we state the undecidability result (as a combination of the statements from [BCLV24, BCV24]) and we prove the analogous result to Theorem 3 stated for Schreier graphs (Theorem 6), which is equivalent to Theorem 3 by Lemma 7. Section 3 is dedicated to the upper approximation of $r$-neighborhood statistics of graphings, and showing the implication mentioned above. In Section 5 we state several open problems. Finally, in Appendix A we present standard encodings between graphs and Schreier graphs and prove Lemma 7.
@@ 2. Proof of Theorem 3
Both sampling regularity for dense graphs and neighborhood statistic regularity are well known to extend to decorated graphs. We will work with graphs equipped with a Schreier decoration of an action of a free group on finitely many generators.
@@
For a rooted edge-labeled graph $F_{\bullet}$ of radius $r$ and an $\mathbb{F}_{d}$ Schreier graph $G$, define the Schreier neighborhood statistic
@@
for all rooted edge-labeled graphs $F_{\bullet}$ of radius $r$.
@@
Theorem 6 (Regularity is Ineffective in Schreier Graphs).
@@
The following theorem summarizes a key step of the argument of Bowen, Chapman, Lubotzky, and Vidick in their resolution of the Aldous–Lyons conjecture. See [BCLV24, Theorem 1.13 and 7.3] and [BCV24, Theorem 1.1] for the general result.
@@ 3. Local statistics of graphings
Since $[0,1]^{\mathcal{F}_{\Delta}(r)}$ is compact in the metric $d_{\infty}$, the image $U_{r}(\mathcal{G}_{\Delta})$ is totally bounded.
@@
for all rooted graphs $F_{\bullet}$ of radius $r$.
@@
[…]
@@
There is a Turing machine $M(d,r,S)$ which takes as input a number of generators $d$, radius $r$, and $S\subseteq[0,1]^{\mathcal{F}^{*}_{d}(r)}$ a union of finitely many open $d_{\infty}$-neighborhoods (centered at rational points) such that $M(d,r,S)$ halts if and only if $U^{*}_{d,r}(\text{IRS}_{d})\subseteq S$.
@@
First we note that the value $U^{*}_{d,r}(\theta)$ is determined by its restriction to $W_{d}(2r)$ (which is a pseudo-IRS). In particular we can extend the map to $U^{*}_{d,r}:\text{P-IRS}_{d}(k)\to[0,1]^{\mathcal{F}^{*}_{d}(r)}$ for any $k\geq 2r$. We have
@@ 4. A Decision Problem for Local Statistics
In light of Remark 11, there is potentially another approach to resolving the Aldous-Lyons conjecture avoiding non-local games by directly showing that for some $\Delta\geq 3$ there is no computable sparse regularity bound $N_{\Delta}(\epsilon,r)$. We point out that this problem can be rephrased into a decision problem about graphs with specified local statistics.
@@
  • if the output is “yes” then there is a finite graph of degree at most $\Delta$ whose $r$-neighborhood statistics is in the set $S$,
  • if the output is “no” then there is no finite graph of degree at most $\Delta$ such that the $\varepsilon$ ball around the $r$-neighborhood statistics is a subset of $S$.
Note that if we allowed $\varepsilon=0$ in the LSDF, then the function would just determine whether there exists a finite graph with specified local statistics. Such a function cannot be computable because the local statistics of a bounded degree graph can encode the run of a Turing machine, such that the machine halts if and only if a finite graph with these local statistics exists. The stronger statement that there is no computable LSDF is in fact equivalent to the non-computability of a sparse regularity bound $N_{\Delta}(\varepsilon,r)$.
@@
Initially we let $\mathscr{G}$ be empty and we recursively apply the following. Let $S$ be the complement in $[0,1]^{\mathcal{F}_{\Delta}(r)}$ of the union of the $\varepsilon/2$ balls around the statistics of each element of $\mathscr{G}$. If $\Phi_{\Delta}(\varepsilon/2,r,S)$ returns “no” then we conclude that $S$ does not contain the $\varepsilon/2$ ball around the statistics of any graph of degree at most $\Delta$. This implies that any graph of degree at most $\Delta$ is within $\varepsilon$ of some graph in $\mathscr{G}$ in $r$-neighborhood statistics, so $\mathscr{G}$ is an $\varepsilon$-net. If $\Phi_{\Delta}(\varepsilon/2,r,S)$ returns “yes” then there is some $G$ of degree at most $\Delta$ with $r$-neighborhood statistics in $S$. By going through a fixed enumeration of finite graphs, we can find such a $G$ with minimal cardinality. We add $G$ to $\mathscr{G}$ and iterate.
@@ Appendix A Encoding graphs and Schreier graphs
$a_{i}$$x$$y$$x_{a_{i}}$$y_{a_{i}^{-1}}$path of length $2d$path of length $i$
Figure 2. Example of encoding of a directed edge
@@
for any labeled rooted graph $F_{\bullet}$ of radius $r$. Therefore we can let $N^{*}_{\Delta}(\varepsilon,r)$ be the maximum number of vertices in one of these decoded graphs $G^{\prime}$. ∎

### v3

@@ 1. Introduction and main results
For sparse graphs, any homomorphism density is close to zero and hence such an approximation is not useful. Instead, one approximates them by neighborhood statistics, defined as follows. Suppose that $G$ is a finite graph with degrees bounded by $\Delta$. A rooted graph of radius $\leq r$ is a graph $F_{\bullet}$ with a distinguished root vertex such that all vertices of $F_{\bullet}$ are of distance $\leq r$ from the root. Rooted graphs are considered up to root-preserving isomorphism $\cong_{\bullet}$ of graphs. For a rooted graph $F_{\bullet}$ of radius $\leq r$ and a (finite) graph $G$ define the corresponding $r$-neighborhood statistic in $G$ as
@@
which is equal to the probability that the radius $r$ ball around a uniformly chosen vertex of $G$ is isomorphic, as a rooted graph, to $F_{\bullet}$.
@@
for all rooted graphs $F_{\bullet}$ of radius $\leq r$.
@@
Bowen, Chapman, Lubotzky, and Vidick employed so-called subgroup tests to encode non-local games and then built on the ideas from the earlier deep results of Ji, Natarajan, Vidick, Wright, and Yuen [JNV+22] to deduce undecidability. We note that although the subgroup test approach might not be necessary for their original goal [Man25] it is indeed essential for our argument. Let us also mention that solely a counterexample to the Aldous–Lyons conjecture in itself would have been likely insufficient to show Theorem 3, but, the insight provided by the undecidability results suffices for this purpose.
@@
In Section 2 we state the undecidability result (as a combination of the statements from [BCLV24, BCV24]) and we prove the analogous result to Theorem 3 stated for Schreier graphs (Theorem 6), which is equivalent to Theorem 3 by Lemma 7. Section 3 is dedicated to the upper approximation of $r$-neighborhood statistics of graphings, and showing the implication mentioned above. In Section 4 we describe a concrete decision function, the inexistence of which would also imply the failure of the Aldous-Lyons conjecture. In Section 5 we state several open problems. Finally, in Appendix A we present standard encodings between graphs and Schreier graphs and prove Lemma 7.
@@ 2. Proof of Theorem 3
Both sampling regularity for dense graphs and neighborhood statistics regularity are well known to extend to decorated graphs. We will work with graphs equipped with a Schreier decoration of an action of a free group on finitely many generators.
@@
For a rooted edge-labeled graph $F_{\bullet}$ of radius $\leq r$ and an $\mathbb{F}_{d}$ Schreier graph $G$, define the Schreier neighborhood statistic
@@
for all rooted edge-labeled graphs $F_{\bullet}$ of radius $\leq r$.
@@
Theorem 6 (Regularity is ineffective for Schreier graphs).
@@
The following theorem summarizes a key step of the argument of Bowen, Chapman, Lubotzky, and Vidick in their resolution of the Aldous–Lyons conjecture. See [BCLV24, Theorems 1.13 and 7.3] and [BCV24, Theorem 1.1] for the general result.
@@ 3. Local statistics of graphings
Since $0,1]^{\mathcal{F}_{\Delta}(r)}$ is compact in the metric $d_{\infty}$, the image $U_{r}(\mathcal{G}_{\Delta})$ is totally bounded.
@@
for all rooted graphs $F_{\bullet}$ of radius $\leq r$.
@@
Given $k>0$ let $W_{d}(k)$ be the ball of radius $k$ around the identity in the Cayley graph of $\mathbb{F}_{d}$ with respect to the standard generating set. A $k$-pseudo-subgroup of $\mathbb{F}_{d}$ is a subset of $W_{d}(k)$ which is the intersection of $W_{d}(k)$ with a subgroup of $\mathbb{F}_{d}$. Then a $k$-pseudo-IRS is defined as a Borel probability measure on $\mathcal{P}(W_{d}(k))\subseteq\mathcal{P}(\mathbb{F}_{d})$ which is invariant under conjugation by the standard generators (when this conjugation is defined). Similarly to the above, we denote the set of $k$-pseudo-IRSs of $\mathbb{F}_{d}$ by $\text{P-IRS}_{d}(k)$. In fact, it follows that
@@
There is a Turing machine $M(d,r,S)$, which takes as input a number of generators $d$, radius $r$, and $S\subseteq[0,1]^{\mathcal{F}^{*}_{d}(r)}$ a union of finitely many open $d_{\infty}$-neighborhoods (centered at rational points) such that $M(d,r,S)$ halts if and only if $U^{*}_{d,r}(\text{IRS}_{d})\subseteq S$.
@@
First, we note that the value $U^{*}_{d,r}(\theta)$ is determined by its restriction to $W_{d}(2r)$ (which is a pseudo-IRS). In particular, we can extend the map to $U^{*}_{d,r}:\text{P-IRS}_{d}(k)\to[0,1]^{\mathcal{F}^{*}_{d}(r)}$ for any $k\geq 2r$. We have
@@ 4. A Decision Problem for Local Statistics
In light of Remark 11, there is a potential alternative approach to resolving the Aldous-Lyons conjecture avoiding non-local games by directly showing that for some $\Delta\geq 3$ there is no computable sparse regularity bound $N_{\Delta}(\epsilon,r)$. We point out that this problem can be rephrased as a decision problem about graphs with specified local statistics.
@@
  • if the output is “yes” then there is a finite graph of degree at most $\Delta$ whose $r$-neighborhood statistic is in the set $S$,
  • if the output is “no” then there is no finite graph of degree at most $\Delta$ such that the $\varepsilon$ ball around the $r$-neighborhood statistic is a subset of $S$.
Note that if we allowed $\varepsilon=0$ in the LSDF, then the function would just determine whether there exists a finite graph with specified local statistics. Such a function cannot be computable since the local statistics of a bounded degree graph can encode the execution of a Turing machine, such that the machine halts if and only if a finite graph with these local statistics exists. The stronger statement that there is no computable LSDF is in fact equivalent to the non-computability of a sparse regularity bound $N_{\Delta}(\varepsilon,r)$.
@@
Initially we let $\mathscr{G}$ be empty and we recursively apply the following. Let $S$ be the complement in $0,1]^{\mathcal{F}_{\Delta}(r)}$ of the union of the $\varepsilon/2$ balls around the statistics of each element of $\mathscr{G}$. If $\Phi_{\Delta}(\varepsilon/2,r,S)$ returns “no” then we conclude that $S$ does not contain the $\varepsilon/2$ ball around the statistics of any graph of degree at most $\Delta$. This implies that any graph of degree at most $\Delta$ is within $\varepsilon$ of some graph in $\mathscr{G}$ in $r$-neighborhood statistics, so $\mathscr{G}$ is an $\varepsilon$-net. If $\Phi_{\Delta}(\varepsilon/2,r,S)$ returns “yes” then there is some $G$ of degree at most $\Delta$ with $r$-neighborhood statistics in $S$. By going through a fixed enumeration of finite graphs, we can find such a $G$ with minimal cardinality. We add $G$ to $\mathscr{G}$ and iterate.
@@ Appendix A Encoding graphs and Schreier graphs
[…]
Figure 2. Example of the encoding of a directed edge
@@
for any labeled rooted graph $F_{\bullet}$ of radius $\leq r$. Therefore we can let $N^{*}_{\Delta}(\varepsilon,r)$ be the maximum number of vertices in one of these decoded graphs $G^{\prime}$. ∎